# A Free-fermion Formulation of Two-dimensional Ising Models

### v1

@@ III.1 The honeycomb lattice
We apply the decorated lattice technique to the honeycomb lattice, as shown in Fig. 4(a). It is easy to verify that, the region bounded by dash lines can be regarded as a vertex site on the square lattice. Each vertex site consists of two honeycomb lattice Ising spins. Fig. 4(b) shows the details with $J_{1}$ determined by Eq. (3). It is natural to introduce a one-to-one mapping from the spin states of $\left\{{\tilde{s}}\right\}$ to the arrow configurations
$\mathop{-\cdot-}~{}\small{+/-}~{}\mapsto~{}\rightarrow/\leftarrow,~{}\mathop{\cdot}\limits_{|}^{|}~{}\small{+/-}~{}\mapsto~{}\uparrow/\downarrow.$ (13)
@@
Figure 4: The diagrammatic representation of applying the decorated lattice technique to the honeycomb lattice model. The region bounded by dash lines in (a) forms a vertex site of the sixteen-vertex model. The details of the vertex unit are shown in (b).
III.1.1 Zero field
In the zero field, the partition function can be expressed by
$Z=A^{N}\sum\limits_{\left\{{\tilde{s}}\right\}=\pm 1}{\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)}$ (14)
@@ III.1.1 Zero field
$\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)=\sum\limits_{{s_{1}}=\pm 1,{s_{2}}=\pm 1}e^{-\beta J{s_{1}}{s_{2}}}\times$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{{e^{-\beta{J_{1}}\left[{{s_{1}}\left({{{\tilde{s}}_{1}}+{{\tilde{s}}_{2}}}\right)+{s_{2}}\left({{{\tilde{s}}_{3}}+{{\tilde{s}}_{4}}}\right)}\right]}}}$ (15)
and $A$ defined in Eq. (4). $\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)$ is the Boltzmann factor of the vertex unit shown in Fig. 4(b). Then the vertex weights can be determined directly from $\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)$, according to the vertex configurations in Fig. 1 and the mapping in Eq. (13). That is
${\omega_{1}}=\omega\left({+,+,+,+}\right)=2{e^{\beta J}}+{e^{-\beta J}}\left({{e^{-4\beta{J_{1}}}}+{e^{4\beta{J_{1}}}}}\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}=2\left({{e^{\beta J}}-{e^{-\beta J}}}\right)+4{e^{-5\beta J}},$
@@
${\omega_{3}}=\omega\left({+,-,-,+}\right)=2\left({{e^{\beta J}}+{e^{-\beta J}}}\right),$
$\omega_{4}=\omega_{5}=\omega_{6}=\omega_{3},$
${\omega_{7}}=\omega\left({-,-,+,+}\right)=2{e^{-\beta J}}+{e^{\beta J}}\left({{e^{-4\beta{J_{1}}}}+{e^{4\beta{J_{1}}}}}\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}=2\left({{e^{-\beta J}}-{e^{\beta J}}}\right)+4{e^{-3\beta J}},$
@@
${\omega_{9}}=\omega\left({+,-,+,+}\right)=\left({{e^{-\beta J}}+{e^{\beta J}}}\right)\left({{e^{-2\beta{J_{1}}}}+{e^{2\beta{J_{1}}}}}\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}=2\left({{e^{-3\beta J}}+{e^{-\beta J}}}\right),$
${\omega_{10}}=\cdots={\omega_{16}}={\omega_{9}}.$ (16)
We see that the resulting sixteen-vertex model is symmetric. As stated before, the vertex weights can be rearranged by performing the weak-graph expansion such that an even eight-vertex model is obtained. The new weights are straightforward to show from Eq. (8)
${{\tilde{\omega}}_{1}}=\frac{1}{2}{\omega_{1}}+{\omega_{3}}+\frac{1}{2}{\omega_{7}}+2{\omega_{9}}=2{e^{\beta J}}{\left({{e^{-2\beta J}}+1}\right)^{3}},$
${{\tilde{\omega}}_{2}}=\frac{1}{2}{\omega_{1}}+{\omega_{3}}+\frac{1}{2}{\omega_{7}}-2{\omega_{9}}$
$~{}~{}~{}~{}=2{e^{\beta J}}\left({{e^{-2\beta J}}+1}\right){\left({{e^{-2\beta J}}-1}\right)^{2}},$
${{\tilde{\omega}}_{3}}={{\tilde{\omega}}_{4}}={{\tilde{\omega}}_{5}}={{\tilde{\omega}}_{6}}={{\tilde{\omega}}_{2}},$
${{\tilde{\omega}}_{7}}=\frac{1}{2}{\omega_{1}}-{\omega_{3}}+\frac{1}{2}{\omega_{7}}=2{e^{\beta J}}{\left({{e^{-2\beta J}}+1}\right)^{2}}\left({{e^{-2\beta J}}-1}\right),$
@@
One can easily examine that the even free-fermion condition Eq. (11) is satisfied. Thus, the honeycomb lattice Ising model in the zero field is equivalent with an even free-fermion model. Denote the number of vertex sites by $N_{\rm{v}}$. The exact solution of the even free-fermion model $\mathop{\lim}\limits_{{N_{{\rm{v}}}}\to\infty}\frac{1}{{{N_{{\rm{v}}}}}}\ln{Z_{{\rm{even}}}}$ had been given in Refs. [41, 42] (note that the ordering of vertices (5)–(8) in Fig. 1 differs slightly from that of Refs. [41, 42], but the expression for the solution is the same). We show the result quoting this well-known solution and noticing Eq. (14)
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z$ $=$ $\ln A+\frac{1}{2}\mathop{\lim}\limits_{{N_{{\rm{v}}}}\to\infty}\frac{1}{{{N_{{\rm{v}}}}}}\ln{Z_{{\rm{even}}}}$
$=$ $\ln\left({\frac{1}{2}{e^{\beta J}}}\right)+\frac{1}{{16{\pi^{2}}}}\int_{0}^{2\pi}d\theta\int_{0}^{2\pi}d\phi\ln\left[a\right.$
$+b\cos\theta$ $+$ $c\cos\phi+d\cos\left({\theta-\phi}\right)+\left.e\left({\theta+\phi}\right)\right],$ (18)
@@
$a$ $=$ ${2^{7}}{e^{-4\beta J}}\left[{{{\cosh}^{3}}\left({2\beta J}\right)+1}\right],$
$b$ $=$ $c={2^{7}}{e^{-4\beta J}}{\sinh^{2}}\left({2\beta J}\right),$
$d$ $=$ $-{2^{7}}{e^{-4\beta J}}{\sinh^{2}}\left({2\beta J}\right),$
$e$ $=$ $0$ (20)
@@
[…]
$~{}~{}~{}~{}\ln\left\{{{\cosh}^{3}}\left({2\beta J}\right)+1\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}\left.+{{\sinh}^{2}}\left({2\beta J}\right)\left[{\cos\theta+\cos\phi+\cos\left({\theta-\phi}\right)}\right]\right\}.$ (21)
It can be easily verified that, this expression is consistent with the results published in the previous literatures [28, 29, 30, 31].
III.1.2 Imaginary field
@@
by using the identity ${i^{{s_{i}}}}=i\times{s_{i}}$ [22]. We can see that, $i^{N}$ is a constant and the product $\prod\limits_{i=1}^{N}{{s_{i}}}$ can be conveniently included in the factors $\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)$. We just need to rewrite the partition function as
$Z=i^{N}A^{N}\sum\limits_{\left\{{\tilde{s}}\right\}=\pm 1}{\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)}$ (23)
@@
$\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)=\sum\limits_{{s_{1}}=\pm 1,{s_{2}}=\pm 1}(s_{1}s_{2})e^{-\beta J{s_{1}}{s_{2}}}\times$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{{e^{-\beta{J_{1}}\left[{{s_{1}}\left({{{\tilde{s}}_{1}}+{{\tilde{s}}_{2}}}\right)+{s_{2}}\left({{{\tilde{s}}_{3}}+{{\tilde{s}}_{4}}}\right)}\right]}}}.$ (24)
@@
$\omega_{1}=\omega_{2}=-2\left({{e^{\beta J}}+{e^{-\beta J}}}\right)+4{e^{-5\beta J}},$
$\omega_{3}=\omega_{4}=\omega_{5}=\omega_{6}=2\left(e^{-\beta J}-e^{\beta J}\right),$
$\omega_{7}=\omega_{8}=2\left({{e^{\beta J}}+{e^{-\beta J}}}\right)-4{e^{-3\beta J}},$
$\omega_{9}=\cdots=\omega_{16}=2\left({{e^{-3\beta J}}-{e^{-\beta J}}}\right).$ (25)
@@
${\tilde{\omega}}_{1}=-16{e^{-2\beta J}}{\cosh^{2}}\left({\beta J}\right)\sinh\left({\beta J}\right),$
${\tilde{\omega}}_{2}=-16{e^{-2\beta J}}{\sinh^{3}}\left({\beta J}\right),$
${{\tilde{\omega}}_{3}}={{\tilde{\omega}}_{4}}={{\tilde{\omega}}_{5}}={{\tilde{\omega}}_{6}}={{\tilde{\omega}}_{1}},$
${{\tilde{\omega}}_{7}}={{\tilde{\omega}}_{8}}=16{e^{-2\beta J}}\cosh\left({\beta J}\right){\sinh^{2}}\left({\beta J}\right).$ (26)
@@
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=i\frac{\pi}{2}+\ln A+\frac{1}{2}\mathop{\lim}\limits_{{N_{{\rm{v}}}}\to\infty}\frac{1}{{{N_{{\rm{v}}}}}}\ln{Z_{{\rm{even}}}}$ (27)

### v2

@@ III.1 The honeycomb lattice
We apply the decorated lattice technique to the honeycomb lattice, as shown in Fig. 4(a). It is easy to verify that, the region bounded by dash lines can be regarded as a vertex site on the square lattice. Each vertex site consists of two honeycomb lattice Ising spins. Fig. 4(b) shows the details with $\tilde{J}_{2}$ and $\tilde{J}_{3}$ determined from $J_{2}$ and $J_{3}$ by Eq. (3), respectively. It is natural to introduce a one-to-one mapping from the spin states of $\left\{{\tilde{s}}\right\}$ in Fig. 4(b) to the arrow configurations
$\text{for horizontal edge}~~\mathop{\frac{~~}{~~}\!\!\bullet\!\!\frac{~~}{~~}}~+/-~\Rightarrow~\rightarrow/\leftarrow,$
$\text{for vertical edge}~~\mathop{\bullet}\limits_{|}^{|}~+/-~\Rightarrow~\uparrow/\downarrow.$ (29)
@@
Figure 4: The diagrammatic representation of applying the decorated lattice technique to the honeycomb lattice model. The region bounded by dashed lines in (a) forms a vertex site of the sixteen-vertex model. The details of the vertex unit are shown in (b).
III.1.1 Zero-field case
In the zero-field case, the partition function can be expressed by
$Z=A_{2}^{N/2}A_{3}^{N/2}\sum\limits_{\left\{{\tilde{s}}\right\}=\pm 1}~{\prod\limits_{{\rm{all~vertices}}}{\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)}}$ (30)
@@ III.1.1 Zero-field case
$\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)=\sum\limits_{{s_{1}}=\pm 1,{s_{2}}=\pm 1}e^{-\beta J_{1}{s_{1}}{s_{2}}}\times$
$~~~~~~~~~~~~~~~~~~~~~~~~~~e^{-\beta\left[{{{\tilde{J}}_{2}}\left({{s_{1}}{{\tilde{s}}_{2}}+{s_{2}}{{\tilde{s}}_{3}}}\right)+{{\tilde{J}}_{3}}\left({{s_{1}}{{\tilde{s}}_{1}}+{s_{2}}{{\tilde{s}}_{4}}}\right)}\right]}$ (31)
and $A_{i}$ determined from $J_{i}$ by Eq. (4). $\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)$ is the Boltzmann factor of the vertex unit shown in Fig. 4(b). Then the vertex weights can be determined directly from $\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)$, according to the vertex configurations in Fig. 1 and the mapping in Eq. (29). That is
$\omega_{1}=\omega\left({++++}\right)=2{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}}_{2}}+{{\tilde{J}}_{3}}}\right)}}+{e^{-2\beta\left({{{\tilde{J}}_{2}}+{{\tilde{J}}_{3}}}\right)}}}\right],$
@@
${\omega_{3}}=\omega\left({+--+}\right)=2{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}}_{2}}-{{\tilde{J}}_{3}}}\right)}}+{e^{-2\beta\left({{{\tilde{J}}_{2}}-{{\tilde{J}}_{3}}}\right)}}}\right],$
$\omega_{4}=\omega_{3},$
${\omega_{5}}=\omega\left({+-+-}\right)=2{e^{-\beta{J_{1}}}}+{e^{\beta{J_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}}_{2}}-{{\tilde{J}}_{3}}}\right)}}+{e^{-2\beta\left({{{\tilde{J}}_{2}}-{{\tilde{J}}_{3}}}\right)}}}\right],$
$\omega_{6}=\omega_{5},$
$\omega_{7}=\omega\left({--++}\right)=2{e^{-\beta{J_{1}}}}+{e^{\beta{J_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}}_{2}}+{{\tilde{J}}_{3}}}\right)}}+{e^{-2\beta\left({{{\tilde{J}}_{2}}+{{\tilde{J}}_{3}}}\right)}}}\right],$
@@
${\omega_{9}}=\omega\left({+-++}\right)$
$~~~~={e^{\beta{J_{1}}}}\left({{e^{2\beta{{\tilde{J}}_{2}}}}+{e^{-2\beta{{\tilde{J}}_{2}}}}}\right)+{e^{-\beta{J_{1}}}}\left({{e^{2\beta{{\tilde{J}}_{3}}}}+{e^{-2\beta{{\tilde{J}}_{3}}}}}\right),$
$\omega_{10}=\omega_{11}=\omega_{12}=\omega_{9},$
$\omega_{13}=\omega\left({-+++}\right)$
$~~~~~={e^{-\beta{J_{1}}}}\left({{e^{2\beta{{\tilde{J}}_{2}}}}+{e^{-2\beta{{\tilde{J}}_{2}}}}}\right)+{e^{\beta{J_{1}}}}\left({{e^{2\beta{{\tilde{J}}_{3}}}}+{e^{-2\beta{{\tilde{J}}_{3}}}}}\right),$
$\omega_{14}=\omega_{15}=\omega_{16}=\omega_{13}.$ (32)
We see that the resulting sixteen-vertex model is symmetric. As stated before, the vertex weights can be rearranged by performing the weak-graph expansion such that an even eight-vertex model is obtained. The new weights are straightforward to show from Eq. (24):
${\tilde{\omega}}_{1}$ $=\frac{1}{2}{\omega_{1}}+\frac{1}{2}{\omega_{3}}+\frac{1}{2}{\omega_{5}}+\frac{1}{2}{\omega_{7}}+{\omega_{9}}+{\omega_{13}}$
$=2\left({{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}}\right)\left({{e^{-2\beta{{J}_{2}}}}+1}\right)\left({{e^{-2\beta{{J}_{3}}}}+1}\right),$
${\tilde{\omega}}_{2}$ $=\frac{1}{2}{\omega_{1}}+\frac{1}{2}{\omega_{3}}+\frac{1}{2}{\omega_{5}}+\frac{1}{2}{\omega_{7}}-{\omega_{9}}-{\omega_{13}}$
$=2\left({{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}}\right)\left({{e^{-2\beta{{J}_{2}}}}-1}\right)\left({{e^{-2\beta{{J}_{3}}}}-1}\right),$
${\tilde{\omega}}_{3}$ $=\frac{1}{2}{\omega_{1}}+\frac{1}{2}{\omega_{3}}-\frac{1}{2}{\omega_{5}}-\frac{1}{2}{\omega_{7}}-{\omega_{9}}+{\omega_{13}}$
$=2\left({{e^{\beta{J_{1}}}}-{e^{-\beta{J_{1}}}}}\right)\left({-{e^{-2\beta{{J}_{2}}}}+1}\right)\left({{e^{-2\beta{{J}_{3}}}}+1}\right),$
${\tilde{\omega}}_{4}$ $=\frac{1}{2}{\omega_{1}}+\frac{1}{2}{\omega_{3}}-\frac{1}{2}{\omega_{5}}-\frac{1}{2}{\omega_{7}}+{\omega_{9}}-{\omega_{13}}$
$=2\left({{e^{\beta{J_{1}}}}-{e^{-\beta{J_{1}}}}}\right)\left({{e^{-2\beta{{J}_{2}}}}+1}\right)\left({-{e^{-2\beta{{J}_{3}}}}+1}\right),$
${\tilde{\omega}}_{5}$ $=\frac{1}{2}{\omega_{1}}-\frac{1}{2}{\omega_{3}}+\frac{1}{2}{\omega_{5}}-\frac{1}{2}{\omega_{7}}$
$=-2\left({{e^{\beta{J_{1}}}}-{e^{-\beta{J_{1}}}}}\right)\left({{e^{-4\beta{{J}_{2}}}}-1}\right)^{1/2}\left({{e^{-4\beta{{J}_{3}}}}-1}\right)^{1/2},$
${\tilde{\omega}}_{6}$ $={\tilde{\omega}}_{5},$
${\tilde{\omega}}_{7}$ $=\frac{1}{2}{\omega_{1}}-\frac{1}{2}{\omega_{3}}-\frac{1}{2}{\omega_{5}}+\frac{1}{2}{\omega_{7}}$
$=2\left({{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}}\right)\left({{e^{-4\beta{{J}_{2}}}}-1}\right)^{1/2}\left({{e^{-4\beta{{J}_{3}}}}-1}\right)^{1/2},$
@@
One can easily examine that the even free-fermion condition Eq. (27) is satisfied. Thus, the honeycomb lattice Ising model in the zero field is equivalent with an even free-fermion model. Denote the number of vertex sites by $N_{\rm{v}}$. The exact solution of the even free-fermion model $\mathop{\lim}\limits_{{N_{{\rm{v}}}}\to\infty}\frac{1}{{{N_{{\rm{v}}}}}}\ln{Z_{{\rm{even}}}}$ had been given in Refs. Fan and Wu (1969, 1970) (note that the ordering of vertices (5)–(8) in Fig. 1 differs slightly from that of Refs. Fan and Wu (1969, 1970), but the expression for the solution is the same). We show the result quoting this well-known solution and noticing Eq. (30)
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=\frac{1}{2}(\ln A_{2}+\ln A_{3})+\frac{1}{2}\mathop{\lim}\limits_{{N_{{\rm{v}}}}\to\infty}\frac{1}{{{N_{{\rm{v}}}}}}\ln{Z_{{\rm{even}}}}$
$=\frac{1}{2}\ln\left({\frac{1}{2}{e^{\beta J_{2}}}}\right)+\frac{1}{2}\ln\left({\frac{1}{2}{e^{\beta J_{3}}}}\right)+\frac{1}{2}\frac{1}{{8{\pi^{2}}}}\int_{0}^{2\pi}d\theta\int_{0}^{2\pi}d\phi$
$~~~~\ln\left[a+b\cos\theta+c\cos\phi+d\cos\left({\theta-\phi}\right)+e\cos\left({\theta+\phi}\right)\right]$ (34)
@@
$a={2^{7}}e^{-2\beta(J_{2}+J_{3})}\left[\cosh\left(2\beta J_{1}\right)\cosh\left(2\beta J_{2}\right)\cosh\left(2\beta J_{3}\right)+1\right],$
$b={2^{7}}e^{-2\beta(J_{2}+J_{3})}\sinh\left(2\beta J_{1}\right)\sinh\left(2\beta J_{2}\right),$
$c={2^{7}}e^{-2\beta(J_{2}+J_{3})}\sinh\left(2\beta J_{1}\right)\sinh\left(2\beta J_{3}\right),$
$d=-{2^{7}}e^{-2\beta(J_{2}+J_{3})}\sinh\left(2\beta J_{2}\right)\sinh\left(2\beta J_{3}\right),$
$e=0.$ (36)
@@
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=\frac{3}{4}\ln 2+\frac{1}{{16{\pi^{2}}}}\int_{0}^{2\pi}d\theta\int_{0}^{2\pi}d\phi$
$~~~~~~~~~~~\ln\left[\cosh\left(2\beta J_{1}\right)\cosh\left(2\beta J_{2}\right)\cosh\left(2\beta J_{3}\right)+1\right$.
$~~~~~~~~~~~~~~~+\sinh\left(2\beta J_{1}\right)\sinh\left(2\beta J_{2}\right)\cos\theta$
$~~~~~~~~~~~~~~~+\sinh\left(2\beta J_{1}\right)\sinh\left(2\beta J_{3}\right)\cos\phi$
$~~~~~~~~~~~~~~~\left.-\sinh\left(2\beta J_{2}\right)\sinh\left(2\beta J_{3}\right)\cos\left(\theta-\phi\right)\right].$ (37)
One can verify that, this expression is consistent with the results published in the previous literatures Husimi and Syôzi (1950); Syôzi (1950); Houtappel (1950); Temperley (1950).
III.1.2 Imaginary-field case
@@
by using the identity ${i^{{s_{i}}}}=i\times{s_{i}}$ Lin and Wu (1988). We can see that, $i^{N}$ is a constant and the product $\prod\limits_{i=1}^{N}{{s_{i}}}$ can be conveniently included in the factors $\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)$. We just need to rewrite the partition function as
$Z=i^{N}A_{2}^{N/2}A_{3}^{N/2}\sum\limits_{\left\{{\tilde{s}}\right\}=\pm 1}~{\prod\limits_{{\rm{all~vertices}}}{\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)}}$ (39)
@@
$\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)=\sum\limits_{{s_{1}}=\pm 1,{s_{2}}=\pm 1}(s_{1}s_{2})e^{-\beta J_{1}{s_{1}}{s_{2}}}\times$
$~~~~~~~~~~~~~~~~~~~~~~~~~{{e^{-\beta\left[{{{\tilde{J}}_{2}}\left({{s_{1}}{{\tilde{s}}_{2}}+{s_{2}}{{\tilde{s}}_{3}}}\right)+{{\tilde{J}}_{3}}\left({{s_{1}}{{\tilde{s}}_{1}}+{s_{2}}{{\tilde{s}}_{4}}}\right)}\right]}}}.$ (40)
@@
$\omega_{1}=\omega_{2}=-2{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}}_{2}}+{{\tilde{J}}_{3}}}\right)}}+{e^{-2\beta\left({{{\tilde{J}}_{2}}+{{\tilde{J}}_{3}}}\right)}}}\right],$
$\omega_{3}=\omega_{4}=-2{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}}_{2}}-{{\tilde{J}}_{3}}}\right)}}+{e^{-2\beta\left({{{\tilde{J}}_{2}}-{{\tilde{J}}_{3}}}\right)}}}\right],$
$\omega_{5}=\omega_{6}=2{e^{-\beta{J_{1}}}}-{e^{\beta{J_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}}_{2}}-{{\tilde{J}}_{3}}}\right)}}+{e^{-2\beta\left({{{\tilde{J}}_{2}}-{{\tilde{J}}_{3}}}\right)}}}\right],$
$\omega_{7}=\omega_{8}=2{e^{-\beta{J_{1}}}}-{e^{\beta{J_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}}_{2}}+{{\tilde{J}}_{3}}}\right)}}+{e^{-2\beta\left({{{\tilde{J}}_{2}}+{{\tilde{J}}_{3}}}\right)}}}\right],$
$\omega_{9}=\omega_{10}=\omega_{11}=\omega_{12}$
$~~~={-e^{\beta{J_{1}}}}\left({{e^{2\beta{{\tilde{J}}_{2}}}}+{e^{-2\beta{{\tilde{J}}_{2}}}}}\right)+{e^{-\beta{J_{1}}}}\left({{e^{2\beta{{\tilde{J}}_{3}}}}+{e^{-2\beta{{\tilde{J}}_{3}}}}}\right),$
$\omega_{13}=\omega_{14}=\omega_{15}=\omega_{16}$
$~~~~~={e^{-\beta{J_{1}}}}\left({{e^{2\beta{{\tilde{J}}_{2}}}}+{e^{-2\beta{{\tilde{J}}_{2}}}}}\right)-{e^{\beta{J_{1}}}}\left({{e^{2\beta{{\tilde{J}}_{3}}}}+{e^{-2\beta{{\tilde{J}}_{3}}}}}\right).$ (41)
@@
${\tilde{\omega}}_{1}$ $=2\left({-{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}}\right)\left({{e^{-2\beta{{J}_{2}}}}+1}\right)\left({{e^{-2\beta{{J}_{3}}}}+1}\right),$
${\tilde{\omega}}_{2}$ $=2\left({-{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}}\right)\left({{e^{-2\beta{{J}_{2}}}}-1}\right)\left({{e^{-2\beta{{J}_{3}}}}-1}\right),$
${\tilde{\omega}}_{3}$ $=-2\left({{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}}\right)\left({-{e^{-2\beta{{J}_{2}}}}+1}\right)\left({{e^{-2\beta{{J}_{3}}}}+1}\right),$
${\tilde{\omega}}_{4}$ $=-2\left({{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}}\right)\left({{e^{-2\beta{{J}_{2}}}}+1}\right)\left({-{e^{-2\beta{{J}_{3}}}}+1}\right),$
${\tilde{\omega}}_{5}$ $={\tilde{\omega}}_{6}$
$=2\left({{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}}\right)\left({{e^{-4\beta{{J}_{2}}}}-1}\right)^{1/2}\left({{e^{-4\beta{{J}_{3}}}}-1}\right)^{1/2},$
${\tilde{\omega}}_{7}$ $={\tilde{\omega}}_{8}$
$=2\left({-{e^{\beta{J_{1}}}}+{e^{-\beta{J_{1}}}}}\right)\left({{e^{-4\beta{{J}_{2}}}}-1}\right)^{1/2}\left({{e^{-4\beta{{J}_{3}}}}-1}\right)^{1/2}.$ (42)
@@
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=i\frac{\pi}{2}+\frac{1}{2}(\ln A_{2}+\ln A_{3})+\frac{1}{2}\mathop{\lim}\limits_{N_{\rm{v}}\to\infty}\frac{1}{N_{\rm{v}}}\ln Z_{\rm{even}}$ (43)
@@
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=i\frac{\pi}{2}+\frac{3}{4}\ln 2+\frac{1}{{16{\pi^{2}}}}\int_{0}^{2\pi}d\theta\int_{0}^{2\pi}d\phi$
$~~~~~~~~~~~\ln\left[\cosh\left(2\beta J_{1}\right)\cosh\left(2\beta J_{2}\right)\cosh\left(2\beta J_{3}\right)-1\right$.
$~~~~~~~~~~~~~~~+\sinh\left(2\beta J_{1}\right)\sinh\left(2\beta J_{2}\right)\cos\theta$
$~~~~~~~~~~~~~~~+\sinh\left(2\beta J_{1}\right)\sinh\left(2\beta J_{3}\right)\cos\phi$
$~~~~~~~~~~~~~~~\left.+\sinh\left(2\beta J_{2}\right)\sinh\left(2\beta J_{3}\right)\cos\left(\theta-\phi\right)\right].$ (44)
One can examine that our result recovers the proposed solutions of this case Baxter (1965); Lin and Wu (1988).
III.2 The triangular lattice
To deal with the triangular lattice Ising model, we employ the star-triangle transformation as shown in Fig. 5(a). The transformed spin system is on the honeycomb lattice, with the replacement of the nearest-neighbour interactions $J_{i}\to J^{\prime}_{i}$. The new anisotropic interactions $J^{\prime}_{i}$ are determined by Eq. (6). Applying the decorated lattice technique in the same way as that in Fig. 4(a), we construct a sixteen-vertex model of which the vertex site is shown Fig. 5(b). The mapping from the spin states to the arrow configurations is still that in Eq. (29). The relation between $J^{\prime}_{i}$ and ${\tilde{J}^{\prime}}_{i}$ is the same with that in Eq. (3).
III.2.1 Zero-field case
In the zero field, the partition function can be obtained using the result for the honeycomb lattice as:
$Z={B^{N}}Z\left({\rm{honeycomb}},J_{i}\to J^{\prime}_{i},H_{\rm{ex}}=0\right)$